# Pro-Pose: Unpaired Full-Body Portrait Synthesis via Canonical UV Maps

### v1

@@ Abstract
Photographs of people taken by professional photographers typically present the person in beautiful lighting, with an interesting pose, and flattering quality. This is unlike common photos people can take of themselves. In this paper, we explore how to create a “professional” version of a person’s photograph, i.e., in a chosen pose, in a simple environment, with good lighting, and standard black top/bottom clothing. A key challenge is to preserve the person’s unique identity, face and body features while transforming the photo. If there would exist a large paired dataset of the same person photographed both “in the wild” and by a professional photographer, the problem would potentially be easier to solve. However, such data does not exist, especially for a large variety of identities. To that end, we propose two key insights: 1) Our method transforms the input photo and person’s face to a canonical UV space, which is further coupled with reposing methodology to model occlusions and novel view synthesis. Operating in UV space allows us to leverage existing unpaired datasets. 2) We personalize the output photo via multi image finetuning. Our approach yields high-quality, reposed portraits and achieves strong qualitative and quantitative performance on real-world imagery.
Figure 1: Pro-Pose Avatar Generation. From a single in-the-wild photo (top row), Pro-Pose generates a portfolio of high-fidelity avatars (bottom rows), fully driven by arbitrary SMPL-X poses [36]. The appearance is canonicalized to a black tank top and shorts and photoshoot setting environment. Our method preserves the user’s identity, facial features, and body shape across novel views and complex poses.
@@ 1 Introduction
Images have become the foundation of digital identity, yet everyday photos are often captured in uncontrolled conditions, with cluttered backgrounds, imperfect lighting, and arbitrary poses and clothing. Consequently, obtaining a clean, controllable “photoshoot-style” representation is difficult, primarily due to the scarcity of high-quality training data. Existing reposing methods such as MCLD [24] and LEFFA [53] rely on the small-scale paired DeepFashion dataset [25] ($\approx$100 identities), leading to overfitting and limited generalization to unseen individuals, and they cannot modify garments. In contrast, virtual try-on (VTO) models [9] typically rely on paired data of a source person and a target garment. Crucially, during training, they derive the person input from the target image and thus are trained to preserve the original pose rather than control or modify it. Consequently, neither method can jointly repose a person and generalize to real-world imagery - motivating our goal to generate photorealistic, reposed portraits of a person in standardized photoshoot conditions that preserve identity while abstracting away clothing complexity. As shown in Figure 1, our approach can produce a versatile portfolio of reposed, fashion-style portraits from a single input.
@@
In this paper, we address the challenge of limited identity diversity by leveraging both abundant single images [16, 56, 55] and scarce paired data [25]. We propose a self-supervised framework utilizing a canonical UV-space [36], which theoretically decouples pose from texture. However, practical extraction remains ill-posed, since the specific boundaries of the visible texture map inevitably leak pose information. To eliminate these clues, we introduce Donor-based UV Reposing. This method uses occlusion masks from unrelated “donor” images to disguise the original visibility patterns, forcing the model to learn robust geometric warping. This formulation enables seamless joint training on 30K paired samples and $\approx$470K unpaired single images. Finally, we employ Gemini 2.5 Flash Image [7] to standardize the garments in the target images, creating a Base Clothing (BC) dataset featuring subjects in black tops and shorts with fixed faces and poses.
@@
Our method employs a dual-branch training strategy to leverage both paired and unpaired data. The paired branch conditions the model on a SMPL-X [36] target, the partial UV texture map, and a source face crop to synthesize the subject in the standardized base garment. Conversely, for unpaired data, we utilize a self-supervised approach where the input also serves as the target. Here, we extract a donor-based UV map and drop out the face condition. This forces the network to reconstruct identity solely from the warped texture, preventing it from trivially copying input pixels. Finally, to address the overfitting risks inherent in single-shot personalization methods trained on limited data, we introduce a lightweight finetuning strategy. This adapts the pre-trained priors to new subjects using minimal reference images, ensuring high identity preservation without requiring large-scale multi-image datasets.
@@
We summarize our contributions as follows: (1) a scalable joint training framework that integrates scarce paired data [25] with abundant unpaired single images [16, 56, 55] to overcome the limited identity diversity of existing benchmarks; (2) a novel self-supervised Donor-based UV Reposing mechanism that reduces coupling between pose and texture to prevent boundary-based pose leakage, enabling effective learning from unpaired data; (3) a stragety to create a large scale Base Clothing (BC[25, 16], synthesized via Gemini 2.5 Flash Image [7] to ensure consistent visual conditions; (4) a subject-specific adaptation mechanism that utilizes our unified training objective to fine-tune LORA layers at test-time, significantly reducing identity drift in challenging scenarios; and (5) extensive evaluations demonstrating that our approach achieves realistic, identity-faithful synthesis on challenging real-world imagery.
@@ 2 Related Work
2D Human Image Synthesis. Research in 2D human image synthesis has historically split into related tracks: pose-guided generation and disentangled attribute editing. Pose-guided methods aim to generate a novel pose of a person from a single image [41, 57, 21, 37]. To this aim, PG2 [29] adopted a coarse-to-fine pipeline, while subsequent single-stage architectures improved end-to-end training by incorporating spatial alignment via deformable skip connections [41], progressive attention [57], or dense appearance flow [21]. Parallel work focused on disentangling attributes, such as appearance and shape [5], or attribute-specific editing [31, 50]. This line of work was further generalized by motion models for arbitrary object animation [42]. To reduce paired data needs, [40] introduced cyclic self-supervision. However, these 2D models often struggle with spatial consistency and entangle pose with texture, leading to artifacts and identity loss.
@@
3D-Aware Human Generation. To overcome spatial inconsistency, later methods incorporated 3D reasoning, implicit surface modeling, or neural rendering [38, 39, 12, 46, 52]. 3DHumanGAN [49] introduced a 3D-pose mapping module to enforce geometric plausibility, and EG3D [2] established efficient hybrid explicit/implicit 3D GANs for view-consistent generation. Other work such as PIFu [38] and PIFuHD [39] recovered pixel-aligned implicit fields, while ConTex-Human [8] focused on achieving texture consistency. More recent approaches have explored 3D Gaussian Splatting for real-time avatars [17, 32, 33] and self-supervised learning from social media videos [13, 14]. However, these approaches rely on strict multi-view or 3D supervision. While such datasets may be voluminous in terms of total frames, they remain severely limited in identity diversity compared to 2D images, preventing generalization to the distribution of real-world human appearances.
@@
Diffusion-Based Animation. Recent diffusion-based techniques have significantly advanced fidelity and controllability [26, 24, 53, 11, 44, 28, 1]. Coarse-to-Fine Latent Diffusion (CFLD) [26] separates semantic and texture modeling; Animate Anyone [11] and DisCo [44] extend this paradigm to temporal consistency and disentangled control; and follow-up works such as MCLD [24] and LEFFA [53] enhance pose controllability. While high-quality, these models remain highly data-hungry, relying on large paired datasets like DeepFashion. Consequently, they are prone to overfitting and “identity drift” on unseen individuals.
@@
Virtual Try-On and Garment Manipulation. Virtual try-on (VTO) methods focus on transferring garments onto target bodies while preserving identity [9, 43, 48, 19, 34, 47, 56, 55]. VITON [9] and CP-VTON [43] pioneered geometric matching for in-shop clothing. M&M VTO [56] extends this field with multi-garment diffusion control. However, these methods assume garment–person pairs rather than person-pose pairs, limiting their ability to handle reposing.
@@ 3 Method
Figure 2: Pose Leakage Mitigation. (a) Standard partial textures $\mathbf{T_{p}}$ leak source pose information via occlusion boundaries, allowing trivial reconstruction shortcuts. (b) Generating pseudo-pairs via image-space rendering is computationally prohibitive for online training. (c) Our Donor-based Reposing efficiently bypasses rendering by applying random donor masks $\mathbf{M}_{\tilde{\mathbf{p}}}$ directly in UV space. This simulates novel occlusions, preventing leakage and forcing the network to learn robust geometric warping.
Figure 3: Overview of our Avatar Generation Framework. Our approach leverages single-view datasets by operating in a canonical UV space, extracting UV texture and pose [36]. Left (Paired Supervision): When ground-truth pose pairs are available, we condition the Flow Matching model directly on the target pose and face crop. Right (Single-View Self-Supervision): To prevent ”pose leakage” from occlusion boundaries when training on single images, we introduce a Donor-based UV Reposing module (Sec. 3.2.1). This synthetically re-poses the input texture using a random donor visibility mask, forcing the model to learn robust identity representations. Furthermore, we drop-out the face crop condition in this branch to prevent trivial reconstruction via pixel-perfect information leakage.
@@ 3.1 Overview and Problem Formulation
The fundamental goal is to generate an image, $\mathbf{A}_{\mathbf{p}}$, of a specific person in a novel target pose $\mathbf{p}$ (represented as a SMPL-X [36] rendering). Given a single reference image $\mathbf{I}_{\mathbf{p}^{\prime}}$ of that person in a source pose $\mathbf{p}^{\prime}$, we formulate this as a conditional generation problem:
@@
A standard approach to train such a model involves minimizing a reconstruction loss, such as $\|\mathbf{A}_{\mathbf{p}}-\mathbf{I}_{\mathbf{p}}\|$, where $\mathbf{I}_{\mathbf{p}}$ is a ground-truth image of the same person in the target pose $\mathbf{p}$. This requires a large dataset of pose-paired images $(\mathbf{I}_{\mathbf{p}^{\prime}},\mathbf{I}_{\mathbf{p}})$. However, collecting such data at scale presents significant challenges due to scarcity, limited identity diversity, and inconsistencies (e.g., clothing changes) between pairs.
These limitations motivate leveraging abundant unpaired single images. However, naively setting $\mathbf{p}=\mathbf{p}^{\prime}$ leads to a trivial identity function. The central challenge is therefore to formulate a self-supervised manner that forces the model to learn meaningful representations from a single image.
@@ 3.2 Self-Supervision in Canonical UV Space
In practice, single-view unwrapping is ill-posed due to self-occlusions. We can only obtain a partial texture map, $\mathbf{T}_{\mathbf{p}}=\mathbf{T}\odot\mathbf{M}_{\mathbf{p}}$, where $\mathbf{M}_{\mathbf{p}}\in\{0,1\}^{H\times W}$ is the binary visibility mask uniquely defined by pose $\mathbf{p}$ (see Figure 2 a). We must therefore approximate the generator using a partial information:
@@
While Eq. 3 enables self-supervision, it introduces a critical challenge: the partial texture $\mathbf{T}_{\mathbf{p}}$ is highly correlated with the pose $\mathbf{p}$ via the mask $\mathbf{M}_{\mathbf{p}}$. As visualized in Figure 2 a, occlusion boundaries in $\mathbf{T}_{\mathbf{p}}$ perfectly align with the target pose. This allows the network to minimize reconstruction loss trivially by ”copy-pasting” visible pixels based on these boundaries, rather than learning the desired 3D geometric warping.
@@ 3.3 Generative Framework
We parameterize our generator $g_{\theta}$ as a Latent Rectified Flow model [23], instantiated using the state-of-the-art Flux.1 [dev] transformer backbone [18]. Consistent with this architecture, we operate in the compressed 16-channel latent space defined by the pre-trained Flux Variational Autoencoder (VAE), comprising an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$.
Target images $\mathbf{I}_{\mathbf{p}}$ are compressed into latent representations $\mathbf{x}_{0}=\mathcal{E}(\mathbf{I}_{\mathbf{p}})$. Following the Flow Matching formulation [23], we model the generative process as an Ordinary Differential Equation (ODE) [3] that transports samples from a standard Gaussian noise distribution $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the data distribution $\mathbf{x}_{0}$.

### v2

@@ Abstract
Photographs of people taken by professional photographers typically present the person in beautiful lighting, with an interesting pose, and flattering quality. This is unlike common photos people take of themselves in uncontrolled conditions. In this paper, we explore how to canonicalize a person’s "in-the-wild" photograph into a controllable, high-fidelity avatar—reposed in a simple environment with standardized minimal clothing. A key challenge is preserving the person’s unique whole-body identity, facial features, and body shape while stripping away the complex occlusions of their original garments. While a large paired dataset of the same person in varied clothing and poses would simplify this, such data does not exist. To that end, we propose two key insights: 1) Our method transforms the input photo into a canonical full-body UV space, which we couple with a novel reposing methodology to model occlusions and synthesize novel views. Operating in UV space allows us to decouple pose from appearance and leverage massive unpaired datasets. 2) We personalize the output photo via multi-image finetuning to ensure robust identity preservation under extreme pose changes. Our approach yields high-quality, reposed portraits that achieve strong quantitative performance on real-world imagery, providing an ideal, clean biometric canvas that significantly improves the fidelity of downstream applications like Virtual Try-On (VTO).
Keywords:
Pose-guided Person Image Synthesis, Virtual Try-On, Avatar Generation
Figure 1: Pro-Pose Canonicalization and Downstream Virtual Try-On. Given a single in-the-wild photo (left), Pro-Pose synthesizes high-fidelity, pose-controllable avatars (middle) driven by arbitrary SMPL-X [Pavlakos2019SMPLX] poses. By canonicalizing the subject into a minimal black outfit, our method preserves the user’s visible identity – facial features, skin textures, and body shape, while stripping away the occlusions of the original clothing. This standardized representation provides an optimal, clean geometric canvas for downstream tasks like Virtual Try-On (right), enabling off-the-shelf VTO [googlevto2023] models to apply novel garments without interference from the source image.
@@ 1 Introduction
Images have become the foundation of digital identity, yet everyday photos are often captured under uncontrolled conditions with cluttered backgrounds, imperfect lighting, and arbitrary poses and clothing. Consequently, obtaining a clean, controllable “photoshoot-style” representation is difficult, primarily due to scarce high-quality training data. Existing reposing methods such as MCLD [MCLD] and LEFFA [Zhou2025LEFFA] rely on the small-scale paired DeepFashion dataset [deepfashion] ($\approx$100 identities), leading to overfitting and limited generalization to unseen individuals, and they cannot modify garments. In contrast, virtual try-on (VTO) models [Han2018VITON] typically rely on paired data of a source person and a target garment. Crucially, during training, they derive the person input from the target image and thus are trained to preserve the original pose rather than control or modify it. Consequently, neither method can jointly repose a person and generalize to real-world imagery - motivating our goal to generate photorealistic, reposed portraits that preserve full-body identity (face, body shape, and skin features) in standardized conditions. Beyond generating standalone digital avatars, this controlled canonical representation acts as a powerful intermediate foundation for downstream tasks like VTO. Current VTO models often suffer from unwanted structural leakage or geometric guidance from the subject’s original clothing, which frequently leads to failures on unconstrained in-the-wild photos. By deliberately canonicalizing the subject into a minimal black tank top and shorts, Pro-Pose provides VTO pipelines with an effective, form-fitting ’blank canvas’ that robustly preserves body shape and pose while eliminating the geometric interference of source garments. As shown in Figure 1, our approach produces a versatile portfolio of reposed, identity-faithful portraits from a single input.
@@
In this paper, we address the limited identity diversity by leveraging both abundant single images [ffhq, Zhu2024MMVTO, Cheng2023TryOnDiffusion] and scarce paired data [deepfashion]. We propose a self-supervised framework utilizing a canonical UV-space [Pavlakos2019SMPLX], which theoretically decouples pose from texture. However, practical extraction remains ill-posed, since the boundaries of the visible texture map inevitably leak pose information. To eliminate these clues, we introduce Donor-based UV Reposing, which uses occlusion masks from unrelated “donor” images to disguise the original visibility patterns, forcing the model to learn robust geometric warping. This formulation enables seamless joint training on 30K paired samples and $\approx$470K unpaired single images. Finally, we employ Gemini 2.5 Flash Image [nanobanana] to standardize the garments in the target images, creating a Base Clothing (BC) dataset featuring subjects in black tops and shorts without altering their identity or pose.
@@
Our method employs a dual-branch training strategy to leverage both paired and unpaired data. The paired branch conditions the model on a SMPL-X [Pavlakos2019SMPLX] target, the partial UV texture map, and a source face crop to synthesize the subject in the standardized base garment. Conversely, for unpaired data, we utilize a self-supervised approach where the input also serves as the target. Here, we extract a donor-based UV map and drop out the face condition. This forces the network to reconstruct identity solely from the warped texture, preventing it from trivially copying input pixels. Finally, to address the overfitting risks inherent in single-shot personalization methods trained on limited data, we introduce a lightweight finetuning strategy. This adapts the pre-trained priors to new subjects using minimal reference images, ensuring high identity preservation without requiring large-scale multi-image datasets.
@@
We summarize our contributions as follows: (1) a scalable joint training framework that integrates scarce paired data [deepfashion] with abundant unpaired single images [ffhq, Zhu2024MMVTO, Cheng2023TryOnDiffusion] to overcome the limited identity diversity of existing benchmarks; (2) a novel self-supervised Donor-based UV Reposing mechanism that reduces coupling between pose and full-body texture to prevent boundary-based pose leakage, enabling effective learning from unpaired data; (3) a strategy to create a large-scale Base Clothing (BC) dataset [deepfashion, ffhq], synthesized via Gemini 2.5 Flash Image [nanobanana] to ensure consistent visual conditions; (4) a subject-specific adaptation mechanism that utilizes our unified training objective to fine-tune LoRA layers at test-time, significantly reducing identity drift in challenging scenarios; and (5) extensive evaluations - including comparisons against Gemini 2.5 Flash Image [nanobanana] and demonstrations of downstream VTO - showing that our approach achieves state-of-the-art, identity-faithful synthesis on challenging real-world imagery.
@@ 2 Related Work
2D Human Image Synthesis. Research in 2D human image synthesis has historically split into related tracks: pose-guided generation and disentangled attribute editing. Pose-guided methods aim to generate a novel pose of a person from a single image [Siarohin2018DefGAN, Zhu2019PATN, Li2019DIAF, Ren2020GFLA]. To this aim, PG2 [Ma2017PG2] adopted a coarse-to-fine pipeline, while subsequent single-stage architectures improved end-to-end training by incorporating spatial alignment via deformable skip connections [Siarohin2018DefGAN], progressive attention [Zhu2019PATN], or dense appearance flow [Li2019DIAF]. Parallel work focused on disentangling attributes, such as appearance and shape [Esser2018VAE], or attribute-specific editing [Men2020ADGAN, Park2021PISE]. This line of work was further generalized by motion models for arbitrary object animation [Siarohin_2019_PAMI]. To reduce paired data needs, [Sanyal2021SPICE] introduced cyclic self-supervision. However, these 2D models struggle with spatial consistency and entangle pose with texture, leading to artifacts and identity loss.
@@
3D-Aware Human Generation. To overcome spatial inconsistency, later methods incorporated 3D reasoning, implicit surface modeling, or neural rendering [Saito2019PIFu, Saito2020PIFuHD, Huang2020ARCH, Xiu2022ICON, Weng2022HumanNeRF]. 3DHumanGAN [Yang2023ThreeDHumanGAN] introduced a 3D-pose mapping module to enforce geometric plausibility, and EG3D [Chan2022EG3D] established efficient hybrid explicit/implicit 3D GANs for view-consistent generation. Other work such as PIFu [Saito2019PIFu] and PIFuHD [Saito2020PIFuHD] recovered pixel-aligned implicit fields, while ConTex-Human [Gao2024ConTexHuman] focused on achieving texture consistency. More recent approaches have explored 3D Gaussian Splatting for real-time avatars [kerbl3Dgaussians, moreau2023human, moreau2024human] and self-supervised learning from social media videos [Jafarian_2022_TPAMI, Jafarian_2021_CVPR_TikTok]. However, these approaches rely on strict multi-view or 3D supervision. While such datasets may be voluminous in terms of total frames, they remain severely limited in identity diversity compared to 2D images, preventing generalization to the distribution of real-world human appearances.
@@
Diffusion-Based Animation. Recent diffusion-based techniques have significantly advanced fidelity and controllability [CFLD, MCLD, Zhou2025LEFFA, Hu2024AnimateAnyone, Bar2023DisCo, Lugmayr_2022_CVPR, brooks2022instructpix2pix]. Coarse-to-Fine Latent Diffusion (CFLD) [CFLD] separates semantic and texture modeling; Animate Anyone [Hu2024AnimateAnyone] and DisCo [Bar2023DisCo] extend this paradigm to temporal consistency and disentangled control; and follow-up works such as MCLD [MCLD] and LEFFA [Zhou2025LEFFA] enhance pose controllability. While high-quality, these models remain highly data-hungry, relying on large paired datasets like DeepFashion. Consequently, they are prone to overfitting and “identity drift” on unseen individuals.
@@
Virtual Try-On and Garment Manipulation. Virtual try-on (VTO) methods transfer garments onto target bodies while preserving identity [Han2018VITON, Wang2018CPVTON, Yang2020ACGPN, Lee2022HRVITON, Morelli2022DressCode, Wang2023OOTDiffusion, Zhu2024MMVTO, Cheng2023TryOnDiffusion]. VITON [Han2018VITON] and CP-VTON [Wang2018CPVTON] pioneered geometric matching for in-shop clothing. M&M VTO [Zhu2024MMVTO] extends this with multi-garment diffusion control. However, these methods assume garment–person pairs rather than person-pose pairs, limiting reposing ability.
@@ 3.1 Overview and Problem Formulation
The fundamental goal is to generate an image, $\mathbf{A}_{\mathbf{p}}$, of a specific person in a novel target pose $\mathbf{p}$ (represented as a SMPL-X [Pavlakos2019SMPLX] rendering). Given a single reference image $\mathbf{I}_{\mathbf{p}^{\prime}}$ of that person in a source pose $\mathbf{p}^{\prime}$, we formulate this as a conditional generation problem:
@@
A standard approach to train such a model involves minimizing a reconstruction loss, such as $\|\mathbf{A}_{\mathbf{p}}-\mathbf{I}_{\mathbf{p}}\|$, where $\mathbf{I}_{\mathbf{p}}$ is a ground-truth image of the same person in the target pose $\mathbf{p}$. This requires a large dataset of pose-paired images $(\mathbf{I}_{\mathbf{p}^{\prime}},\mathbf{I}_{\mathbf{p}})$. However, collecting such data at scale presents significant challenges due to scarcity, limited identity diversity, and inconsistencies (e.g. clothing changes) between pairs.
These limitations motivate leveraging abundant unpaired single images. However, naively setting $\mathbf{p}=\mathbf{p}^{\prime}$ leads to a trivial identity function. The central challenge is therefore to formulate self-supervision that forces the model to learn meaningful representations from a single image.
@@ 3.2 Self-Supervision in Canonical UV Space
In practice, single-view unwrapping is ill-posed due to self-occlusions. We can only obtain a partial texture map, $\mathbf{T}_{\mathbf{p}}=\mathbf{T}\odot\mathbf{M}_{\mathbf{p}}$, where $\mathbf{M}_{\mathbf{p}}\in\{0,1\}^{H\times W}$ is the binary visibility mask uniquely defined by pose $\mathbf{p}$ (see Figure 2 a). We must therefore approximate the generator using partial information:
@@
While Eq. 3 enables self-supervision, it introduces a critical challenge: the partial texture $\mathbf{T}_{\mathbf{p}}$ is highly correlated with the pose $\mathbf{p}$ via the mask $\mathbf{M}_{\mathbf{p}}$. As visualized in Figure 2 a, occlusion boundaries in $\mathbf{T}_{\mathbf{p}}$ perfectly align with the target pose. This allows the network to minimize reconstruction loss trivially by "copy-pasting" visible pixels based on these boundaries, rather than learning the desired 3D geometric warping.
@@ 3.2.1 Donor-Based UV Reposing
Figure 2: Pose Leakage Mitigation. (a) Standard partial textures $\mathbf{T_{p}}$ leak source pose information via occlusion boundaries, allowing trivial reconstruction shortcuts. (b) Generating pseudo-pairs via image-space rendering is computationally prohibitive for online training. Moreover, SMPL-X lacks fine details—such as hair—leading to unrealistic renderings. (c) Our Donor-based Reposing efficiently bypasses rendering by applying random donor masks $\mathbf{M}_{\tilde{\mathbf{p}}}$ directly in UV space. This simulates novel occlusions, preventing leakage and forcing the network to learn robust geometric warping. Note: To clearly visualize the subtle boundary differences (indicated by arrows), here we display a crop of just the face region from the full texture map; in practice, donor-reposing applies to all body parts.
@@ 3.3 Generative Framework
We parameterize our generator $g_{\theta}$ as a Latent Rectified Flow model [lipman2023flowmatchinggenerativemodeling], instantiated using the state-of-the-art Flux.1 [dev] transformer backbone [flux2024]. Consistent with this architecture, we operate in the compressed 16-channel latent space defined by the pre-trained Flux Variational Autoencoder (VAE), comprising an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$.
Target images $\mathbf{I}_{\mathbf{p}}$ are compressed into latent representations $\mathbf{x}_{0}=\mathcal{E}(\mathbf{I}_{\mathbf{p}})$. Following the Flow Matching formulation [lipman2023flowmatchinggenerativemodeling], we model the generative process as an Ordinary Differential Equation (ODE) [NeuralODE] that transports samples from a standard Gaussian noise distribution $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the data distribution $\mathbf{x}_{0}$.